Dorothy is 6 years older than Ricardo. The product of their present ages is twice what the product of their ages was 6 years ago. how old is Dorothy?
I have trued to find something similar in the section of word problem and just totally confused myself- HELP PLEASE!

Found 2 solutions by user_dude2008, stanbon:
Answer by user_dude2008(1862)

"Dorothy is 6 years older than Ricardo." -----> d=r+6

"The product of their present ages is twice what the product of their ages was 6 years ago" ----> d*r=2(d-6)(r-6)


d*r=2(d-6)(r-6)

(r+6)*r=2(r+6-6)(r-6)

(r+6)*r=2r(r-6)

r+6=2(r-6)

r+6=2r-12

6+12=2r-r

18=r

r=18 ----> Ricardo is 18 yrs old


d=r+6=18+6=24 ----> Dorothy is 24 yrs old

Answer by stanbon(75887)

Dorothy is 6 years older than Ricardo. The product of their present ages is twice what the product of their ages was 6 years ago. how old is Dorothy?
------------------
NOW DATA:
Ricardo's age is x ; Dorothy's age is x+6
----------------------------------------
6 Yrs. ago DATA:
Ricardo was x-6 ; Dorothy was x
---------------
Equation:
x(x+6) = 2(x(x-6))
x^2 + 6x = 2(x^2-6x)
x^2 + 6x = 2x^2 -12x
x^2 -18x = 0
x(x-18) = 0
x = 0 or x = 18
-------------------
Realistic answer: x = 18 (Ricardo's age NOW)
x+6 = 24 (Dorothy's age NOW)
===================================
Cheers,
Stan H.